OpenStudy (anonymous):

What is the height of cylinder with a surface area of 226.08 square meters and a radius of 3 meters? (Use 3.14 for π.)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Surface Area of a Cylinder = 2 π r\(^2\) + 2 π r h r = radius h = height Can you solve for "h" from here?

OpenStudy (anonymous):

Surface area of a cylinder: \[A=2\pi r^2+2\pi rh\] Plug in the values: \[226.08=2(3.14)(3)^2+2(3.14)(3)h\] Simplifying the equation: \[226.08=56.52+18.84h\] Subtract 56.52: \[169.56=18.84h\] Divide by 18.84: h=9